# Towards Next-Generation Urban Decision Support Systems through AI-Powered Generation of Scientific Ontology using Large Language Models - A Case in Optimizing Intermodal Freight Transportation

### v1

@@ 2.6 Challenge and Knowledge Gap
Despite the presence of well-established methodologies, ontology management software, and advanced web semantic languages for designing, developing, and implementing scientific ontologies, the creation of efficient and reusable ontologies to support smart city services frequently encounters both methodological and technical challenges. These challenges can impede the effective promotion and integration of such ontologies within the framework of smart city initiatives.
Methodologically, creating scenario-based ontologies tailored to enhance data discovery, integration, and management for specific smart city services, such as parking availability, building control optimization, and traffic volume prediction, is a crucial but repetitive and labor-intensive task in the development of numerous urban decision support tools. This effort becomes increasingly challenging as new data sources and urban simulation scenarios continue to emerge rapidly across expansive urban regions, along with the deployment of new sensors and IoT-connected infrastructure, as well as the increasing data availability from various sources (e.g., crowd-sourcing, autonomous vehicles, and simulation outputs). Moreover, ensuring the reusability and interoperability of newly created scenario-based ontologies remains a significant challenge. For instance, integrating the ontology of a new multimodal freight transportation system with that of an existing urban transportation system can pose complex challenges. The real-time traffic sensor data described in the urban transportation system ontology is also crucial for supporting the operations of the multimodal freight transport and logistics system.
@@
Technically, developing a scenario-based ontology for newly emerging data sources or simulation models often requires time-consuming and labor-intensive efforts. This process demands in-depth domain expertise to extract and interpret essential knowledge from extensive documentation, research articles, and manuals—knowledge that is critical for the specific scope of a smart city application. After extracting the core knowledge, which is subsequently represented as high-level concepts, properties, and relationships, maintaining the semantic consistency of the derived ontology across existing domain ontologies is crucial. This ensures that new data and simulations can be interpreted correctly and uniformly, and readily integrated into the existing software tools. However, this task requires complex and iterative processes, including the standardization of data and metadata, semantic mapping, as well as comprehensive validation and testing.
Given the exceptional capabilities of LLMs in natural language understanding, pattern recognition, reasoning, and self-learning, there is a compelling need to develop an AI-powered autonomous tool for ontology and knowledge graph generation. This tool would leverage LLMs’ capabilities to automate the extraction of knowledge from urban research articles and technical manuals, thereby facilitating the construction of scenario-based ontologies. Such a tool would enhance data-driven research and benefit smart city applications by optimizing complex urban management decisions.
@@ 3 Methodology
This section begins with an overview of the background and design requirements for our proposed ontology generation workflow. It then details the development of our methodology, both conceptually and technically. The ontology produced through our approach is validated by comparing it with the well-known Pizza Ontology [^1] Available at: https://drive.google.com/file/d/1PMM-G9KOKrmJwzrqoYnC-VLnAT1fEgZG/view?usp=sharing, which is frequently used in tutorials for popular ontology software.
Figure 2: A subset of sample data from the three datasets combined to demonstrate the origin-destination pairs, as well as different modes of transportation and types of transit across the United States.
3.1 Design Requirement
Our initiative to prototype an LLM-powered autonomous agent for ontology generation is driven by the need to create a large-scale, data-driven digital twin designed to support complex decision-making for optimizing the intermodal transportation system across the contiguous US. Our digital twin, debuted as ARPA-E RECOIL, leverages the latest state-of-the-art intermodal transportation datasets and simulation tools to optimize the freight transport system. This comprehensive approach takes into account supply and demand dynamics, delivery times, and costs, along with potential socioeconomic impacts such as traffic congestion, air pollution, and carbon emissions. Consequently, it necessitates the creation of an ontology behind the proposed urban digital twin to integrate multi-domain datasets and facilitate the coupling simulations across a variety of disciplines. The target users of the proposed autonomous agents include urban researchers, operational scientists, and software developers with domain expertise and either entry-level or intermediate computer science and programming skills. Detailed design requirements of our autonomous agent is as follows:
@@ 3.1 Design Requirement
    Compatibility and Portability: The autonomous agent should be designed as a generalized workflow, implementable in popular programming languages like Python. It should be modular and flexible, enabling deployment and execution on widely-used online platforms such as Jupyter Notebook.
@@
    Automated Knowledge Source Acquisition: The autonomous agent should be integrated with online databases and archives for scientific research literature via APIs. This integration will enable the agent to automatically download relevant research papers and technical manuals based on user-defined keywords related to a specific scope, dataset, or simulation model.
@@
    Autonomous Knowledge Extraction: The autonomous agent should be capable of comprehending human natural language, recognizing domain-specific terminologies, such as identifying concepts as entities and categories as classes, and understanding the relationships between various concepts and classes.
@@
    Automated Ontology Design: The autonomous agent should adhere to the practices and procedures of well-established ontology design methodologies, including Methontology, OTKM, and NeON (see Section 2.2). This adherence will enable it to create robust knowledge representations characterized by high granularity, reusability, flexibility, and semantic consistency.
@@
    Ontology Interoperability: The autonomous agent should possess the capability to export ontologies as knowledge graphs or in widely accepted ontology languages, such as OWL and RDF (see Section 2.4). This functionality will enable the ontologies to be interpreted by popular ontology management software tools, facilitating validation processes and the creation of database schema for data modeling and integration.
@@
    Self-validating: The autonomous agent should have the capability to validate the ontologies it generates by referencing existing domain ontologies or relevant literature. This process should include checking the definitions of concepts, relationships, and semantic mappings to ensure accuracy and consistency.
The autonomous agent, designed and developed based on the aforementioned requirements, will later be employed to create a sample ontology as a proof-of-concept. This will involve integrating emerging datasets and simulation outputs from the intermodal transportation sector, including the Freight Analysis Framework (FAF), Freight Transportation Optimization Tool (FTOT), and Bureau of Transportation Statistics (BTS). This integration aims to facilitate informed and comprehensive decisions to optimize complex intermodal freight transportation systems. More details regarding the data and simulation used in this work are provided in Section 4.1.
@@ 3.2 Methodological Foundation
The conceptual design of our autonomous agent is depicted in Figure 3, which comprises four primary modules: (a) Knowledge Source Acquisition, (b) Knowledge Source Pre-processing, (c) LLM-Powered Ontology Generation, and (d) Ontology Implementation. Each component encompasses a series of procedures implemented using Python libraries. Additionally, the entire process is hosted in a Python JupyterLab environment to facilitate human supervision of the workflow step-wise. Detailed workflow and procedures within individual modules are presented through the following subsections.
Figure 3: Overall architecture of the proposed autonomous agent for ontology generation with four major modules
@@ 3.2.1 Knowledge Source Acquisition and Pre-Processing
The knowledge source acquisition module automates the discovery, downloading, and archiving of digitized files (e.g., PDFs and XMLs). These files contain essential domain knowledge and are retrieved based on user-defined keywords. These keywords may refer to specific urban or environmental datasets, scientific simulations, and more. The process utilizes SerpApi, a web scraper that identifies and downloads internet resources (e.g., the HTML page of a report or article). Additionally, it employs the Scopus API, an academic database that discovers and downloads research articles using these keywords as search criteria. This API-powered component searches the internet for resources that match the search criteria and downloads them. The resources, including research articles and technical manuals for datasets or simulation models, are saved in common formats such as PDF, HTML, and XML. They are then stored in a designated file directorate, which can be accessed by other components of our autonomous agent.
@@
The knowledge source pre-Processing module is designed to prepare digital files, previously downloaded, into formats suitable for input into the LLM for ontology generation. This module integrates content analysis and data extraction libraries with custom Python scripts to initially cleanse the digital files. This cleansing process removes unnecessary components such as headers, footers, page numbers, and special characters, which could potentially impair the performance of subsequent natural language processing tasks and the LLM’s reasoning capabilities. Subsequently, the content is extracted and transformed into plain text (for text blocks and tables) or images (for figures), depending on their original format. The extracted content is then divided into subsets that comply with the word count limitations of the LLM for each request, specifically using the ChatGPT API in this study.
3.2.2 LLM-powered Ontology Generation
@@
The LLM-powered ontology generation module leverages ChatGPT’s natural language processing (NLP) and reasoning capabilities to automatically generate scenario-based ontologies. This process is facilitated through the use of theory-guided prompts, which are developed based on the core principles of the Methontology approach.
Before initiating the ontology development procedures, plain text and image inputs are processed through two separate workflows to enhance input data quality. Text inputs undergo a series of NLP procedures including tokenization, stop word removal, sentence boundary detection, and text normalization. Concurrently, images are processed through the GPT image recognition engine via API to extract embedded text and summarize the content in text form.
@@
The refined text inputs from both plain texts and images are then fed into the ChatGPT-4 API, which serves as the central reasoning engine, to construct the ontology. The Methontology framework’s procedures have been translated into structured prompts that utilize the flexibility of Python scripting and the ChatGPT API, creating a maintainable workflow. This workflow instructs the LLM on how to create scientific ontologies based on pre-processed input data.
Detailed data flow and its procedure for processing the plain text inputs are described as the following:
@@
    Sentence Boundary Detection: Following text normalization, this process identifies the start and end points of sentences within the text. Accurate sentence boundary detection is essential for understanding the text’s structure and for facilitating subsequent per-sentence processing. It segments a large text into manageable, analyzable units. This functionality is also integrated into the LLM through the ChatGPT API.
@@
    Image Text Extraction: The process uses Optical Character Recognition (OCR) technology to detect and convert text within images into machine-readable characters. In academic figures, this includes extracting annotations, axis labels, data points, and descriptive text, essential for interpreting the graphical content. This process is conducted outside the ChatGPT API using the Pytesseract library.
@@
    GPT Image Recognition Engine: a component of a larger AI system that utilizes models trained to recognize and interpret content within images. This engine processes images to identify objects, features, and patterns, often using advanced neural networks optimized for image analysis. In the context of research articles, this could involve recognizing graphical data, text within figures, or even the structure of the figures themselves. This functionality is integrated into the LLM through the ChatGPT API.
@@
    Summary of Images as Texts: The process synthesizes key information from an image into a concise textual summary. It uses AI to interpret recognized objects and text, distilling these elements to highlight the essential insights or data presented. This is especially useful in academic settings for quickly conveying the significance and results depicted in complex figures or diagrams. This functionality is also integrated into the LLM through the ChatGPT API.

### v3

@@ 3 Knowledge Gaps and Contributions
Despite the availability of well-established methodologies, ontology management software, and advanced web semantic languages for designing, developing, and implementing scientific ontologies, enabling automated creation of efficient and reusable ontologies to support smart city services often encounters challenges and technical limitations.
3.1 Domain Application Challenges
In the field of smart city management, creating scenario-based ontologies is crucial for enhancing data discovery, integration, and management across various urban sectors, such as parking availability, building energy optimization, and traffic volume prediction. This is a sophisticated task that often requires large volumes and diverse types of data, as well as outputs from simulations and predictive analytics. The efforts involved in data preparation and provisioning are also essential for formulating informed decision support. However, these tasks are typically repetitive, time-consuming, and demand specialized domain expertise.
As new data sources and urban simulation scenarios rapidly emerge across expansive urban regions, and with the deployment of new sensors and IoT-connected infrastructure, the challenge intensifies. Additionally, the increasing availability of data from various sources, such as crowd-sourcing, autonomous vehicles, and simulation outputs, further complicates the landscape. From a data interoperability perspective, ontologies and knowledge graphs play critical roles in multi-domain data integration and simulation coupling (Xu et al., 2020; Chen et al., 2018). For instance, solving complex optimization problems in freight transportation systems often requires integrating data on freight, fuel and emission simulations, real-time traffic, and transportation infrastructure. These comprehensive tasks can definitely benefit from an ontological approach. Developing system-based knowledge representation using ontologies and knowledge graphs is essential for discovering, managing, and sharing complex urban science data, which are indispensable for creating urban digital twins, information systems, and decision support tools.
@@ 3.1 Domain Application Challenges
Despite the critical importance of urban knowledge representation and semantic techniques in the smart city sector, designing and constructing scenario-based ontologies for newly emerging data sources or simulation models often require significant time and specialized effort. This process demands deep domain expertise to extract and interpret essential knowledge from extensive documentation, research articles, and manuals. After extracting core knowledge and representing it as high-level concepts, properties, and relationships, it is vital to maintain the semantic consistency of the derived ontology with existing domain ontologies. This ensures that new data and simulations are interpreted correctly, uniformly, and can be seamlessly integrated into existing data and software tools. However, this task involves complex and iterative processes, including data and metadata standardization, semantic mapping, and thorough validation and testing. Consequently, there is a pressing need for automated methods to design and construct ontologies and knowledge graphs that describe newly emerging urban data and simulation outputs.
3.2 Methodological Challenges
There has been a recent surge in research interest in developing more efficient and automated methods for ontology and knowledge construction by leveraging cutting-edge LLMs. However, existing studies in this area are still rare and in their preliminary stages. Our review identified only four relevant studies, including a few ongoing research efforts presented through preprints.
Among these, two contemporary studies focus on proving the feasibility of using LLMs to construct ontologies and knowledge representations from text inputs using zero-shot prompting methods and general-purpose queries (Babaei Giglou et al., 2023; Caufield et al., 2024). These studies provide important proof-of-concepts, demonstrating the potential of LLMs with simple text documents. However, they do not focus on training or guiding LLMs with standard, system-based knowledge engineering methodologies to tackle complex ontology construction tasks involving sophisticated scientific data, which encompasses quantitative fields and multiple data features (such as spatial, temporal, and topological dimensions). The proper interpretation of logic and entity relationships among multiple quantitative fields in scientific data and simulation outputs has not been addressed in previous studies, often requiring LLMs to follow specialized knowledge engineering methodologies. One study explored the potential for automating the creation of domain ontologies and knowledge graphs using text inputs, LLMs, and RAG through the pre-definition of competency questions (Kommineni et al., 2024). This approach aims to formulate a more self-organizing method for creating foundational domain knowledge using diverse text content. However, the research does not focus on creating scenario-based ontologies and knowledge graphs based on specific data sources and simulation tools with particular application objectives for supporting operations science. Another study proposed a novel entity, domain, and context (EDC) method to improve the extraction of entity-relation triplets from generic real-world text documents and automate the generation of knowledge graphs (Zhang and Soh, 2024). This study implemented its proposed method using multiple LLM frameworks and compared their performances. Despite its unique contributions, the study does not emphasize constructing ontological knowledge representation in a hierarchical structure but instead focuses on encoding knowledge into a graph structure. Additionally, the study is not designed to extract knowledge and information from domain-specific text inputs that describe scientific datasets and simulation tools. As a result, it does not emphasize domain-specific relation extraction, definition, or knowledge fusion with existing domain ontologies.
3.3 Knowledge Gaps
Based on our literature review, we have identified several gaps in the existing studies on LLM-ontology. The following aspects have been rarely addressed in previous research.
Complex Domain applications in Operations Science:
Many previous studies have concentrated on testing the feasibility of constructing ontologies and knowledge using LLMs through proof-of-concept and pilot studies. However, there is a gap in utilizing AI-generated ontologies to support real-world decision-making and evaluating their performance in practical applications.
Building Concepts with Complex Features:
Most previous studies have utilized generic text inputs to test the ontology and knowledge graph creation capabilities of LLMs. However, these studies did not involve the handling and interpretation of complex scientific datasets with various data dimensions and features, such as topological, spatial, and temporal aspects.
Integration of Knowledge Engineering Methods:
Many previous studies have employed general queries with zero-shot prompting methods to generate knowledge representation using LLMs. However, the aspect of training or instructing LLMs using commonly-accepted knowledge engineering methodologies, as discussed in Section 2.3, has been rarely addressed.
Interoperability of LLM-generated Content:
While previous studies have focused on testing the feasibility and accuracy of LLM-generated knowledge representations, they often overlook the aspects of interoperability and usability, particularly concerning their application to practical research problems. Potential application areas include supporting the formulation of decision support strategies, guiding the design of critical software and database components, and facilitating the construction of knowledge bases for practical purposes.
Based on the findings and insights from previous studies, we believe there is a need to test the LLM-ontology capabilities on more complex scientific datasets, simulations, and their manuals. This will help create AI-powered methodologies that support knowledge extraction and management, ultimately addressing complex problems in operations science.
3.4 Motivation and Contributions
Building on prior research into the use of Large Language Models (LLMs) for constructing ontologies and knowledge graphs, we aim to advance the current state-of-the-art by exploring LLMs’ capabilities to identify, extract, and interpret concepts in more sophisticated scientific datasets and simulation outputs, which are often quantitative and multi-dimensional. Our goal is to create knowledge representations with practical benefits to support complex decision-making in operations research.
Based on this motivation, we present the design, implementation, and application of an integrated methodology that utilizes the well-accepted METHONTOLOGY framework to instruct an LLM in a few-shot learning fashion. This approach aims to construct scenario-based ontologies from urban research articles, technical manuals, and data subsets. An integral part of our methodology includes a workflow that provides options for utilizing the AI-generated knowledge representations to (a) guide the design and development of critical software and database components for urban decision support systems and (b) facilitate the formulation of decision support strategies, which involve the definition decision metrics and workflows. Due to the interdisciplinary nature of our study, our research contributes to the fields of operations science, software engineering, and the applications of LLMs.
Contributions to LLM Application:
We explored the ability of Large Language Models (LLMs) to understand and interpret entities and relationships in complex spatial and topological datasets and simulation outputs, guided by the METHONTOLOGY framework. Our approach is demonstrated through Section 5.2.
Contributions to Software Engineering:
We demonstrate a practice of utilizing AI-generated ontology to guide software design and semi-automatically generate critical components of a data provisioning system essential for developing an urban decision support system. Our approach aims to systematically bridge AI-generated knowledge representation with existing ontology management tools and database technologies (detaied in Section 5.3), supporting practical tasks in the development of decision support software.
Contributions to Operations Science:
We devised an automated AI-powered workflow, driven by integrated data, simulations, and knowledge representations, to develop intelligent and advanced capabilities for supporting decision-making and optimizing complex urban systems. A detailed demonstration is presented through Section 5.5.
@@ 4 Methodology
This section begins with an overview of the background and design requirements for our proposed ontology generation workflow. It then details the development of our methodology, both conceptually and technically. The ontology produced through our approach is validated by comparing it with the well-known Pizza Ontology, which is frequently used in tutorials for popular ontology software.
4.1 Design Requirements
Our initiative to prototype an LLM-powered workflow for ontology generation is driven by the need to create a large-scale, data-driven digital twin designed to support complex decision-making for optimizing the intermodal transportation system across the contiguous US. Our digital twin, debuted as ARPA-E RECOIL, leverages the latest state-of-the-art intermodal transportation datasets and simulation tools to optimize the freight transport system. This comprehensive approach takes into account supply and demand dynamics, delivery times, and costs, along with potential socioeconomic impacts such as traffic congestion, air pollution, and carbon emissions. Consequently, it necessitates the creation of an ontology behind the proposed urban digital twin to integrate multi-domain datasets and facilitate the coupling simulations across a variety of disciplines. The target users of the proposed automated workflow include urban researchers, operational scientists, and software developers with domain expertise and either entry-level or intermediate computer science and programming skills. Detailed design requirements of our automated workflow are as follows:
@@ 4.1 Design Requirements
    Compatibility and Portability: The workflow should be designed as a generalized workflow, implementable in popular programming languages like Python. It should be modular and flexible, enabling deployment and execution on widely-used online platforms such as Jupyter Notebook.
@@
    Automated Knowledge Source Acquisition: The workflow should be integrated with online databases and archives for scientific research literature via APIs. This integration will enable the agent to automatically download relevant research papers and technical manuals based on user-defined keywords related to a specific scope, dataset, or simulation model.
@@
    Autonomous Knowledge Extraction: The workflow should be capable of comprehending human natural language, recognizing domain-specific terminologies, such as identifying concepts as entities and categories as classes, and understanding the relationships between various concepts and classes.
@@
    Automated Ontology Design: The workflow should adhere to the practices and procedures of well-established ontology design methodologies, including Methontology, OTKM, and NeON (see Section 2.2). This adherence will enable it to create robust knowledge representations characterized by high granularity, reusability, flexibility, and semantic consistency.
@@
    Ontology Interoperability: The workflow should be able to export ontologies as knowledge graphs or in widely accepted ontology languages, such as OWL and RDF (see Section 2.4). This functionality will enable the ontologies to be interpreted by popular ontology management software tools, facilitating validation processes and the creation of database schema for data modeling and integration.
@@
    Self-validating: The autonomous tool should have the capability to validate the ontologies it generates by referencing existing domain ontologies or relevant literature. This process should include checking the definitions of concepts, relationships, and semantic mappings to ensure accuracy and consistency.
Our proposed method, designed and developed based on the aforementioned requirements, will later be employed to create a sample scientific ontology as a proof-of-concept. This will involve integrating emerging datasets and simulation outputs from the intermodal transportation sector, including the Freight Analysis Framework (FAF) and Freight Transportation Optimization Tool (FTOT). This integration aims to facilitate informed and comprehensive decisions to optimize complex intermodal freight transportation systems. More details regarding the data and simulation used in this work are provided in Section 5.1.
@@ 4.2 Methodological Foundation
The conceptual design of our autonomous workflow is depicted in Figure 2, which comprises four primary modules: (a) Knowledge Source Acquisition, (b) Knowledge Source Pre-processing, (c) LLM-Powered Ontology Generation, and (d) Ontology Implementation. Each component encompasses a series of procedures implemented using Python libraries. Additionally, the entire process is based in Python, combining JupyterLab and the use of scripting for automated orchestrated execution of the process. This facilitates human supervision of the workflow step-wise. Detailed workflow and procedures within individual modules are presented through the following subsections.
Figure 2: Overall architecture of the applied workflow for ontology generation with four major modules
@@ 4.2.1 Knowledge Source Acquisition and Pre-Processing
The knowledge source acquisition module automates the discovery, downloading, and archiving of digitized files (e.g., PDFs and XMLs). These files contain essential domain knowledge and are retrieved based on user-defined keywords. These keywords may refer to specific urban or environmental datasets, scientific simulations, and more. It employs the Scopus API, an academic database that discovers and downloads research articles using these keywords as search criteria. This API-powered component searches the internet for resources that match the search criteria and downloads them. The resources, including research articles and technical manuals for datasets or simulation models, are saved in formats like PDF, HTML, and XML. They are then stored in a designated file directorate, which other components of our automated workflow can access.
@@
The knowledge source pre-processing module is designed to prepare digital files, previously downloaded, into formats suitable for input into the LLM for ontology generation. This module integrates content analysis and data extraction libraries with custom Python scripts to initially cleanse the digital files. This cleansing process removes unnecessary components such as headers, footers, page numbers, and special characters, which could potentially impair the performance of subsequent natural language processing tasks and the LLM’s. Subsequently, the content is extracted and transformed into plain text (for text blocks and tables) or images (for figures), depending on their original format. The extracted content is then divided into subsets that comply with the word count limitations of the LLM for each request, specifically using the ChatGPT API in this study.
4.2.2 LLM-powered Ontology Construction - Input Preparation
@@
The LLM-powered ontology generation module leverages ChatGPT’s natural language processing (NLP) and reasoning capabilities to automatically generate scenario-based ontologies. This process is facilitated through the use of prompt templates, which are developed based on the core principles of the Methontology framework.
@@
Before initiating the ontology development procedures, plain text and image inputs are processed through two separate workflows to enhance input data quality. Text inputs undergo a series of NLP procedures, including tokenization, stop word removal, sentence boundary detection, and text normalization. Concurrently, images are processed through the GPT image recognition engine via API to extract embedded text and summarize the content in text form. The refined text inputs from both plain texts and images are then fed into the ChatGPT-4 API, which serves as the central reasoning engine, to construct the ontology. The Methontology framework’s procedures have been translated into structured prompts that utilize the flexibility of Python scripting and the ChatGPT API, creating a maintainable workflow. This workflow instructs the LLM on how to create scientific ontologies based on pre-processed input data. Detailed data flow and its procedure for processing the plain text inputs are described as the following:
@@
    Sentence Boundary Detection: Following text normalization, this process identifies the start and end points of sentences within the text. Accurate sentence boundary detection is essential for understanding the text’s structure and for facilitating subsequent per-sentence processing. It segments a large text into manageable, analyzable units. This functionality is also incorporated into the LLM via the ChatGPT API.
@@
    Image Context Recognition: The process uses Optical Character Recognition (OCR) technology to detect and convert text surrounding images into machine-readable characters. This includes extracting annotations, axis labels, data points, and descriptive text essential for interpreting the graphical content. This process is conducted outside the ChatGPT API using Layout Parser Shen et al. (2021) and the Detectron2 model Wu et al. (2019) for image separation serves as component identification to filter those that represent figures for further processing.
@@
    GPT Image Recognition Engine: a component of a larger AI system that utilizes models trained to recognize and interpret content within images. This engine processes images to identify objects, features, and patterns, often using advanced neural networks optimized for image analysis. In the context of research articles, this involves recognizing graphical data, text within figures, or even the structure of the figures themselves. This functionality is integrated into the LLM through the ChatGPT API.
@@
    Summary of Images as Texts: The process synthesizes key information from an image into a concise textual summary. It uses AI to interpret recognized objects, distilling these elements to highlight the essential insights or data presented. This is especially useful in academic settings for quickly conveying the significance and results depicted in complex figures or diagrams. This functionality is also integrated into the LLM through the ChatGPT API.
4.2.3 LLM-powered Ontology Construction - Knowledge Extraction